### v1

# Optimal Convergence Rate for Mirror Descent Methods with Time-Varying Step Sizes Rules

Abstract.
In this paper, the optimal convergence rate $O\left(N^{-1/2}\right)$ (where $N$ is the total number of iterations performed by the algorithm), without the presence of a logarithmic factor, is proved for mirror descent algorithms with time-varying step sizes, for solving both classical constrained non-smooth problems and problems with the composite model. The proven result is an improvement on the well-known rate $O\left(\log(N)N^{-1/2}\right)$ for the mirror descent algorithms with time-varying step sizes.
Key words and phrases:
Convex Optimization, non-smooth problem, composite problem, mirror descent, mirror-c descent, optimal convergence rate, sub-optimal convergence rate, time-varying step size.
Preprint under update.
1. Introduction
The mirror descent method which originated in [3, 4] and was later analyzed in [5], is considered as the non-Euclidean extension of standard subgradient methods. This method is used in many applications, see [10, 11, 12, 13, 14] and references therein. The standard subgradient methods employ the Euclidean distance function with a suitable step-size in the projection step. Mirror descent extends the standard projected subgradient methods by employing a nonlinear distance function with an optimal step-size in the nonlinear projection step [15]. The mirror descent method is also applicable to optimization problems in Banach spaces where gradient descent is not [16]. An extension of the mirror descent method for constrained problems was proposed in [17, 4]. Usually, the step size and stopping criterion for the mirror descent method require knowing the Lipschitz constant of the objective function and constraints, if any. Adaptive step sizes, which do not require this information, are considered for unconstrained problems in [18], and for constrained problems in [17].
As we pointed out the mirror descent algorithm was proposed and analyzed a long time ago for different step size rules. However, the optimal convergence rate (namely $O(N^{-1/2})$, where $N$ is the total number of iterations performed by the algorithm) of the algorithm was proved only with the fixed (constant) step size. When the step sizes are time-varying (non-constant in all iterations) it proved the sub-optimal convergence rate, namely $O(\log(N)N^{-1/2})$, see [1, 2, 8, 7] and references therein.
The constant step size rule is relatively easy to analyze but has the disadvantage of requiring the a priori knowledge of the total number of iterations employed by the method. In practical situations, the number of iterations is not fixed a priori, and stopping criteria different than merely fixing the total number of iterations is usually imposed. This is why dynamic, or time-varying (i.e., non-constant) step size rules are important [7].
Recently in [9], it was proved the optimal convergence rate for the classical projected subgradient method for time-varying step size. This convergence rate remains the same (optimal) even if we slightly increase the weight of the most recent points, thereby relaxing the ergodic sense.
In this paper, we extend the results of [9] to the mirror descent methods. We proved the optimal convergence rate, without the presence of a logarithmic factor, for mirror descent methods with time-varying step sizes, for solving both classical constrained non-smooth problems (see Sect. 3) and composite problems (see subsection 4).
The paper consists of an introduction and 4 main sections. In Sect. 2 we mentioned the basic facts and tools and mirror descent basics. Sect. 3 devoted to the proof of the optimal convergence rate of the mirror descent algorithm with time-varying step size rules for solving the classical unconstrained non-smooth optimization problem. In Sect. 4 the same analysis, that in Sect. 3, performed for more general optimization problems, namely the composite optimization problem. In Sect. 5 we present some numerical experiments which demonstrate the effectiveness of the proposed analysis.
2. Fundamentals
[…]
$\|y\|_{*}=\max\limits_{x\in\mathbf{E}}\{\langle y,x\rangle:\|x\|\leq 1\},$
where $\langle y,x\rangle$ is the value of the continuous linear functional $y\in\mathbf{E}^{*}$ at $x\in\mathbf{E}$.
Let $Q\subset\mathbf{E}^{n}$ be a closed convex set, and $\psi:Q\longrightarrow\mathbb{R}$ be a proper closed differentiable and $\sigma$-strongly convex (called prox-function or distance generating function). The corresponding Bregman divergence is defined as
$V_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(x),y-x\rangle,\quad\forall x,y\in Q.$
For the Bregmann divergence, it holds the following inequality
(2.1) $V_{\psi}(x,y)\geq\frac{\sigma}{2}\|y-x\|^{2},\quad\forall x,y\in Q.$
The following identity, known as the three points identity, is essential in the analysis of the mirror descent method.
Lemma 2.1.
(three points identity)[6] Suppose that $\psi:\mathbf{E}\longrightarrow(-\infty,\infty]$ is proper closed and convex, and differentiable over $\operatorname{dom}(\partial(\psi))$. Let $a,b\in\operatorname{dom}(\partial(\psi))$ and $c\in\operatorname{dom}(\psi)$. Then it holds
(2.2) $\langle\nabla\psi(b)-\nabla\psi(a),c-a\rangle=V_{\psi}(c,a)+V_{\psi}(a,b)-V_{\psi}(c,b).$
Fenchel-Young inequality. For any $a\in\mathbf{E},b\in\mathbf{E}^{*}$, it holds the following inequality
(2.3) $\left|\langle a,b\rangle\right|\leq\frac{\|a\|^{2}}{2\lambda}+\frac{\lambda\|b\|_{*}}{2},\quad\forall\lambda>0.$
3. Mirror descent algorithm
In this section, let us consider the following optimization problem
(3.1) $\min_{x\in Q}f(x),$
where $f$ is a convex Lipschitz continuous function (non-smooth) with Lipschitz constant $M_{f}>0$, i.e.,
(3.2) $|f(x)-f(y)|\leq M_{f}\|x-y\|,\quad\forall x,y\in Q.$
Here and what follows, we denote the subdifferential of $f$ at $x$ by $\partial f(x)$, and the subgradient of $f$ at any point $x$ by $\nabla f(x)\in\partial f(x)$.
Algorithm 1 Mirror descent method.
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$, number of iterations $N$.
1:  for $k=1,2,\ldots,N$ do
2:     Calculate $\nabla f(x^{k})\in\partial f(x^{k})$,
3:     $x^{k+1}=\arg\min_{x\in Q}\left\{\langle x,\nabla f(x^{k})\rangle+\frac{1}{\gamma_{k}}V_{\psi}(x,x^{k})\right\}$.
4:  end for
Theorem 3.1.
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with a positive non-increasing sequence of step sizes $\{\gamma_{k}\}_{k\geq 1}$, for any fixed $m\geq-1$, it satisfies the following inequality
(3.3) $f(\hat{x})-f(x^{*})\leq\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\left(\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}\right),$
where $\hat{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$, and $x^{*}$ is an optimal solution of (3.1).
Proof.
Let $\tilde{f}(x):=\langle x,\nabla f(x^{k})\rangle+\frac{1}{\gamma_{k}}V_{\psi}(x,x^{k})$, then from Algorithm 1 we have $x^{k+1}=\arg\min\limits_{x\in Q}\tilde{f}(x)$. By the optimality condition, we get
$\langle\nabla\tilde{f}(x^{k+1}),x-x^{k+1}\rangle\geq 0,\quad\forall x\in Q.$
Thus,
$\langle\gamma_{k}\nabla f(x^{k})+\nabla\psi(x^{k+1})-\nabla\psi(x^{k}),x-x^{k+1}\rangle\geq 0,\quad\forall x\in Q.$
In particular for $x=x^{*}$, we have
(3.4) $\langle\nabla\psi(x^{k})-\nabla\psi(x^{k+1})-\gamma_{k}\nabla f(x^{k}),x^{*}-x^{k+1}\rangle\leq 0.$
By using the definition of a subgradient of $f$ at $x^{*}$, we get
$\gamma_{k}\left(f(x^{k})-f(x^{*})\right)$ $\leq\gamma_{k}\langle\nabla f(x^{k}),x^{k}-x^{*}\rangle$
$=\underbrace{\langle\nabla\psi(x^{k})-\nabla\psi(x^{k+1})-\gamma_{k}\nabla f(x^{k}),x^{*}-x^{k+1}\rangle}_{:=J_{1}}$
$\;\;\;\;+\underbrace{\langle\nabla\psi(x^{k+1})-\nabla\psi(x^{k}),x^{*}-x^{k+1}\rangle}_{:=J_{2}}$
$\;\;\;\;+\underbrace{\langle\gamma_{k}\nabla f(x^{k}),x^{k}-x^{k+1}\rangle}_{:=J_{3}}.$
From (3.4), we find $J_{1}\leq 0$. By Lemma 2.1, we have
$J_{2}$ $=-\langle\nabla\psi(x^{k})-\nabla\psi(x^{k+1}),x^{*}-x^{k+1}\rangle$
$=-\left(V_{\psi}(x^{*},x^{k+1})+V_{\psi}(x^{k+1},x^{k})-V_{\psi}(x^{*},x^{k})\right)$
$=V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})-V_{\psi}(x^{k+1},x^{k}).$
By the Fenchel-Young inequality (2.3), with $\lambda=\sigma>0$, we find
$J_{3}\leq\frac{\gamma_{k}^{2}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}+\frac{\sigma}{2}\|x^{k}-x^{k+1}\|^{2}.$
Therefore, we get
$\gamma_{k}\left(f(x^{k})-f(x^{*})\right)$ $\leq V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})-V_{\psi}(x^{k+1},x^{k})$
$\;\;\;\;+\frac{\gamma_{k}^{2}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}+\frac{\sigma}{2}\|x^{k}-x^{k+1}\|^{2}.$
From (2.1), we have
$V_{\psi}(x^{k+1},x^{k})\geq\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}.$
Thus, we get the following inequality
(3.5) $f(x^{k})-f(x^{*})\leq\frac{1}{\gamma_{k}}\left(V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})\right)+\frac{\gamma_{k}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}.$
By multiplying both sides of (3.5) by $\frac{1}{\gamma_{k}^{m}}$, and taking the sum from $1$ to $N$, we get
$\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})-f(x^{*})\right)$ $\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m+1}}\left(V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})\right)$
$\;\;\;\;+\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{2\sigma\gamma_{k}^{m-1}}.$
Since $f$ is a convex function, we have
$\left(\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\right)\left[f\left(\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}\right)-f(x^{*})\right]\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})-f(x^{*})\right).$
Therefore, by setting $\hat{x}:=\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$, we get
$\left(\sum_{k=1}^{N}\gamma_{k}^{-m}\right)\left(f(\hat{x})-f(x^{*})\right)\leq\frac{1}{\gamma_{1}^{m+1}}\left(V_{\psi}(x^{*},x^{1})-V_{\psi}(x^{*},x^{2})\right)$
$\;\;\;\;+\sum_{k=2}^{N-1}\frac{1}{\gamma_{k}^{m+1}}\left(V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})\right)$
$\;\;\;\;+\frac{1}{\gamma_{N}^{m+1}}\left(V_{\psi}(x^{*},x^{N})-V_{\psi}(x^{*},x^{N+1})\right)+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$\leq\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+\frac{1}{\gamma_{N}^{m+1}}V_{\psi}(x^{*},x^{N})+\sum_{k=2}^{N-1}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k})$
$\;\;\;\;-\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{2})-\sum_{k=2}^{N-1}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k+1})+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$=\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k})-\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{2})$
$\;\;\;\;-\sum_{k=3}^{N}\frac{1}{\gamma_{k-1}^{m+1}}V_{\psi}(x^{*},x^{k})+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$=\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k})-\sum_{k=2}^{N}\frac{1}{\gamma_{k-1}^{m+1}}V_{\psi}(x^{*},x^{k})$
$\;\;\;\;+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$\leq\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+V_{\psi}(x^{*},x^{1})\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m+1}}-\frac{1}{\gamma_{k-1}^{m+1}}\right)+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$=\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+V_{\psi}(x^{*},x^{1})\left(-\frac{1}{\gamma_{1}^{m+1}}+\frac{1}{\gamma_{N}^{m+1}}\right)+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
$=\frac{1}{\gamma_{N}^{m+1}}V_{\psi}(x^{*},x^{1})+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Finally, by dividing by $\sum_{k=1}^{N}\gamma_{k}^{-m}$, we get the desired inequality
$f(\hat{x})-f(x^{*})\leq\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\left(\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}\right).$
As a special case of Theorem 3.1, we can deduce the well-known sub-optimal convergence rate of mirror descent algorithm $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, see [1, 2, 8, 7], with the following (adaptive and non-adaptive) time-varying step size rules
(3.6) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\quad\text{or}\quad\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\quad k=1,2,\ldots,N.$
Corollary 3.2.
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with the time-varying step sizes given in (3.6), it satisfies the following sub-optimal convergence rate
(3.7) $f(\tilde{x})-f(x^{*})\leq\frac{M_{f}\left(V_{\psi}(x^{*},x^{1})+1+\log(N)\right)}{\sqrt{\sigma}}\frac{1}{\sqrt{N}}=O\left(\frac{\log(N)}{\sqrt{N}}\right),$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
Proof.
By setting $m=-1$ in (3.3), we get the following inequality
(3.8) $f(\tilde{x})-f(x^{*})\leq\frac{V_{\psi}(x^{*},x^{1})+\frac{1}{2\sigma}\sum_{k=1}^{N}\gamma_{k}^{2}\|\nabla f(x^{k})\|_{*}^{2}}{\sum_{k=1}^{N}\gamma_{k}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (3.8) we find
$f(\tilde{x})-f(x^{*})\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}.$
But
(3.9) $\sum_{k=1}^{N}\frac{1}{k}\leq 1+\log(N),\quad\text{and}\quad\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\geq 2\sqrt{N+1}-2,\quad\forall N\geq 1.$
Therefore,
$f(\tilde{x})-f(x^{*})$ $\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{V_{\psi}(x^{*},x^{1})+1+\log(N)}{2\sqrt{N+1}-2}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}},\quad\forall N\geq 1.$
Where in the last inequality, we used the fact $2\sqrt{2}(\sqrt{N+1}-1)\geq\sqrt{N},\,\forall N\geq 1$.
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (3.8) we find
$f(\tilde{x})-f(x^{*})$ $\leq\frac{V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}=\frac{M_{f}}{\sqrt{2\sigma}}\frac{V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}}.$
Also, from Theorem 3.1 we can obtain the optimal convergence rate of mirror descent algorithm $O\left(\frac{1}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.6).
Corollary 3.3.
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with the time-varying step sizes given in (3.6), it satisfies the following optimal convergence rate
(3.10) $f(\overline{x})-f(x^{*})\leq\frac{M_{f}\left(2+V_{\psi}(x^{*},x^{1})\right)}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$
where $\overline{x}=\frac{1}{N}\sum_{k=1}^{N}x^{k}$.
Proof.
By setting $m=0$ in (3.3), we get the following inequality
(3.11) $f(\overline{x})-f(x^{*})\leq\frac{1}{N}\left(\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\gamma_{k}\|\nabla f(x^{k})\|_{*}^{2}\right),$
where $\overline{x}=\frac{1}{N}\sum_{k=1}^{N}x^{k}$.
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (3.11) we find
$f(\overline{x})-f(x^{*})\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right).$
But
(3.12) $\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\leq 2\sqrt{N},\quad\forall N\geq 1.$
Therefore,
$f(\overline{x})-f(x^{*})\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}V_{\psi}(x^{*},x^{1})+2\sqrt{N}\right)=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (3.11) we find
$f(\overline{x})-f(x^{*})$ $\leq\frac{1}{N}\left(\frac{\sqrt{N}\|\nabla f(x^{N})\|_{*}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{\sqrt{k}}\|\nabla f(x^{k})\|_{*}\right)$
$\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$
$\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}V_{\psi}(x^{*},x^{1})+2\sqrt{N}\right)=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
Remark 3.4.
By setting any $m>-1$ in (3.3) (in Theorem 3.1), the convergence rate of Algorithm 1 with the time-varying step sizes given in (3.6), will be optimal $O\left(\frac{1}{\sqrt{N}}\right)$ without presence of a $\log(N)$ factor. In comparison with the sub-optimal convergence rate (3.7), when $m>0$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points.
4. Mirror descent for composite problem
In this section, we will consider a more general problem than problem (3.1)
(4.1) $\min_{x\in Q}\left\{F(x)=f(x)+h(x)\right\},$
where $Q=\operatorname{dom}(h)\cap\operatorname{dom}(\partial\psi)$ and $f,h$ satisfy the following assumption
Assumption 4.1.
(properties of $f$ and $g$)
  1. (1)
    $f$ and $h$ are proper closed and convex.
  2. (2)
    $\operatorname{dom}(h)\subseteq\operatorname{int}(\operatorname{dom}(f))$.
  3. (3)
    $f$ is $M_{f}$-Lipschitz continuous on $\operatorname{dom}(h)$ for any $M_{f}>0$.
We also, assume that the optimal set of the problem (4.1) is nonempty.
For the problem (4.1), with an additional assumption that $h$ is non-negative, we mention the following well-defined algorithm [1].
Algorithm 2 Mirror-C descent method.
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$, number of iterations $N$.
1:  for $k=1,2,\ldots,N$ do
2:     Calculate $\nabla f(x^{k})\in\partial f(x^{k})$,
3:     $x^{k+1}=\arg\min_{x\in Q}\left\{\gamma_{k}\langle x,\nabla f(x^{k})\rangle+\gamma_{k}h(x)+V_{\psi}(x,x^{k})\right\}$.
4:  end for
Let us mention the following key lemma for the analysis of the Algorithm 2.
Lemma 4.2.
[1] Let
  1. (1)
    $\psi:\mathbf{E}\longrightarrow(-\infty,\infty]$ be a proper closed and convex function differentiable over $\operatorname{dom}(\partial\psi)$,
  2. (2)
    $\varphi:\mathbf{E}\longrightarrow(-\infty,\infty]$ be a proper closed and convex function satisfying $\operatorname{dom}(\partial\varphi)\subseteq\operatorname{dom}(\partial\psi)$,
  3. (3)
    $\psi+\mathbb{I}_{\operatorname{dom}(\varphi)}$ be $\sigma$-strongly convex, where $\mathbb{I}_{A}$ is the indicator function of the set $A$.
Assume that $b\in\operatorname{dom}(\partial(\psi))$, and let $a$ be defined by
$a=\arg\min_{x\in\mathbf{E}}\left\{\varphi(x)+V_{\psi}(x,b)\right\}.$
Then $a\in\operatorname{dom}(\partial(\psi))$, we have
(4.2) $\left\langle\nabla\psi(b)-\nabla\psi(a),u-a\right\rangle\leq\varphi(u)-\varphi(a),\quad\forall u\in\operatorname{dom}(\varphi).$
Theorem 4.3.
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with a positive non-increasing sequence of step sizes $\{\gamma_{k}\}_{k\geq 1}$, for any fixed $m\geq-1$, it satisfies the following inequality
(4.3) $F(\hat{x})-F(x^{*})\leq\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\left(\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}\right),$
where $\hat{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ and $x^{*}$ is an optimal solution of (4.1).
Proof.
From (4.2), with $b=x^{k},a=x^{k+1}$ and $\varphi(x)=\gamma_{k}\langle x,\nabla f(x^{k})\rangle+\gamma_{k}h(x)$, we get the following inequality
(4.4) $\langle\nabla\psi(x^{k})-\nabla\psi(x^{k+1}),u-x^{k+1}\rangle\leq\gamma_{k}\langle u-x^{k+1},\nabla f(x^{k})\rangle+\gamma_{k}h(u)-\gamma_{k}h(x^{k+1}).$
Now, from Lemma three points identity (2.2), with $a=x^{k+1},b=x^{k}$ and $c=u$, we get
(4.5) $\langle\nabla\psi(x^{k})-\nabla\psi(x^{k+1}),u-x^{k+1}\rangle=V_{\psi}(u,x^{k+1})+V_{\psi}(x^{k+1},x^{k})-V_{\psi}(u,x^{k}).$
By combining (4.4) and (4.5), we get the following inequality
(4.6) $V_{\psi}(u,x^{k+1})+V_{\psi}(x^{k+1},x^{k})-V_{\psi}(u,x^{k})$ $\leq\gamma_{k}\langle u-x^{k+1},\nabla f(x^{k})\rangle$
$\;\;\;\;+\gamma_{k}h(u)-\gamma_{k}h(x^{k+1}).$
From (4.6), we find
(4.7) $\gamma_{k}\langle\nabla f(x^{k}),x^{k}-u\rangle$ $+\gamma_{k}h(x^{k+1})-\gamma_{k}h(u)\leq V_{\psi}(u,x^{k})-V_{\psi}(u,x^{k+1})$
$\qquad\quad-V_{\psi}(x^{k+1},x^{k})+\langle\gamma_{k}\nabla f(x^{k}),x^{k}-x^{k+1}\rangle.$
Therefore, by (2.1) and Fenchel-Young inequality (2.3) with $\lambda=\sigma$, from (4.7) we get the following
(4.8) $\gamma_{k}\langle\nabla f(x^{k}),x^{k}-u\rangle$ $+\gamma_{k}h(x^{k+1})-\gamma_{k}h(u)\leq V_{\psi}(u,x^{k})-V_{\psi}(u,x^{k+1})$
$-\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{\gamma_{k}^{2}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}.$
$=V_{\psi}(u,x^{k})-V_{\psi}(u,x^{k+1})+\frac{\gamma_{k}^{2}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}.$
Since $f$ is a convex function, it holds
$f(u)-f(x^{k})\geq\langle\nabla f(x^{k}),u-x^{k}\rangle\Longrightarrow f(x^{k})-f(u)\leq\langle\nabla f(x^{k}),x^{k}-u\rangle,$
and by setting $u=x^{*}$ in (4.8), we get the following inequality
$\gamma_{k}\left(f(x^{k})-f(x^{*})\right)+\gamma_{k}h(x^{k+1})-\gamma_{k}h(x^{*})$ $\leq V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})$
$\;\;\;\;+\frac{\gamma_{k}^{2}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2},$
which means
(4.9) $f(x^{k})+h(x^{k+1})-F(x^{*})\leq\frac{1}{\gamma_{k}}\left(V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})\right)+\frac{\gamma_{k}}{2\sigma}\|\nabla f(x^{k})\|_{*}^{2}.$
By multiplying both sides of (4.9) by $\frac{1}{\gamma_{k}^{m}}$, and taking the sum form $1$ to $N$, we get
$\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})+h(x^{k+1})-F(x^{*})\right)$ $\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m+1}}\left(V_{\psi}(x^{*},x^{k})-V_{\psi}(x^{*},x^{k+1})\right)$
$\;\;\;\;+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Now, in a similar way as in the proof of Theorem 3.1, we conclude that
(4.10) $\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})+h(x^{k+1})-F(x^{*})\right)\leq\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
By adding the term $\frac{h(x^{1})}{\gamma_{1}^{m}}-\frac{h(x^{N+1})}{\gamma_{N}^{m}}$ to both sides of (4.10), we get
the following
$\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})+h(x^{k+1})-F(x^{*})\right)+\frac{f(x^{1})}{\gamma_{1}^{m}}+\frac{h(x^{2})}{\gamma_{1}^{m}}-\frac{F(x^{*})}{\gamma_{1}^{m}}+\frac{h(x^{1})}{\gamma_{1}^{m}}$
$\;\;\;\;-\frac{h(x^{N+1})}{\gamma_{N}^{m}}\leq\frac{h(x^{1})}{\gamma_{1}^{m}}-\frac{h(x^{N+1})}{\gamma_{N}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}},$
Because of $h$ is a non-negative, we have $\frac{h(x^{N+1})}{\gamma_{N}^{m}}>0$, and then we get the following
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})+\frac{1}{\gamma_{k}^{m}}h(x^{k+1})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)+\frac{h(x^{2})}{\gamma_{1}^{m}}$
$\;\;\;\;-\frac{h(x^{N+1})}{\gamma_{N}^{m}}\leq\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}},$
which means
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m}}h(x^{k+1})+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)$
$\;\;\;\;+\frac{h(x^{2})}{\gamma_{1}^{m}}-\frac{h(x^{N+1})}{\gamma_{N}^{m}}\leq\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Thus,
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\sum_{k=3}^{N+1}\frac{1}{\gamma_{k-1}^{m}}h(x^{k})+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)$
$\;\;\;\;+\frac{h(x^{2})}{\gamma_{1}^{m}}-\frac{h(x^{N+1})}{\gamma_{N}^{m}}\leq\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Therefore,
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)+\sum_{k=2}^{N}\frac{1}{\gamma_{k-1}^{m}}h(x^{k})\leq$
$\;\;\;\;\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Thus, we get
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})+\frac{1}{\gamma_{k-1}^{m}}h(x^{k})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)\leq$
$\;\;\;\;\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
Because of the sequence $\{\gamma_{k}\}_{k\geq 1}$ non-increasing, we have $\frac{1}{\gamma_{k-1}^{m}}\geq\frac{1}{\gamma_{k}^{m}}$, and then we conclude
$\sum_{k=2}^{N}\left(\frac{1}{\gamma_{k}^{m}}f(x^{k})+\frac{1}{\gamma_{k}^{m}}h(x^{k})-\frac{1}{\gamma_{k}^{m}}F(x^{*})\right)+\frac{1}{\gamma_{1}^{m}}\left(F(x^{1})-F(x^{*})\right)\leq$
$\;\;\;\;\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}},$
which means
$\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(F(x^{k})-F(x^{*})\right)\leq\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
From the convexity of the function $F$, we have
$\left(\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\right)\left[F\left(\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}\right)-F(x^{*})\right]\leq$
$\;\;\;\;\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(F(x^{k})-F(x^{*})\right).$
Therefore, by setting $\hat{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$, we get the following inequality
(4.11) $\left(\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\right)\left(F(\hat{x})-F(x^{*})\right)\leq\frac{h(x^{1})}{\gamma_{1}^{m}}+\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}^{m+1}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}.$
By dividing both sides of (4.11) by $\sum_{k=1}^{N}\gamma_{k}^{-m}$, we get the desired inequality (4.3).
As a special case of Theorem 4.3, we can deduce the well-known sub-optimal convergence rate of mirror-C descent algorithm 2, namely $O\left(\frac{\log(N)}{N}\right)$, with the time-varying step sizes given in (3.6) (see [1, 2, 8, 7]).
Corollary 4.4.
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with the time-varying step sizes given in (3.6), it satisfies the following sub-optimal convergence rate
(4.12) $F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{\sigma}}\left(\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)\right)\frac{1}{\sqrt{N}}$
$=O\left(\frac{\log(N)}{\sqrt{N}}\right),$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
Proof.
By setting $m=-1$ in (4.3), we get the following inequality
(4.13) $F(\tilde{x})-F(x^{*})\leq\frac{\gamma_{1}h(x^{1})+V_{\psi}(x^{*},x^{1})+\frac{1}{2\sigma}\sum_{k=1}^{N}\gamma_{k}^{2}\|\nabla f(x^{k})\|_{*}^{2}}{\sum_{k=1}^{N}\gamma_{k}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
Case 1 (non-adaptive rule).
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13) we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{2\sqrt{N+1}-2}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}}.$
Where in the last inequality, we used the fact $2\sqrt{2}(\sqrt{N+1}-1)\geq\sqrt{N},\,\forall N\geq 1$.
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.3) we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}}.$
Also, from Theorem 4.3 we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 2, with the time-varying step sizes given in (3.6).
Corollary 4.5.
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with the time-varying step sizes given in (3.6), it satisfies the following optimal convergence rate
(4.14) $f(\overline{x})-f(x^{*})\leq\left(\frac{M_{f}V_{\psi}(x^{*},x^{1})+2M_{f}}{\sqrt{2\sigma}}\right)\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$
where $\overline{x}=\frac{1}{N}\sum_{k=1}^{N}x^{k}$.
Proof.
By setting $m=0$ in (4.3), and since $h(x^{1})=0$, we get the following inequality
(4.15) $F(\overline{x})-F(x^{*})\leq\frac{1}{N}\left(\frac{V_{\psi}(x^{*},x^{1})}{\gamma_{N}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\gamma_{k}\|\nabla f(x^{k})\|_{*}^{2}\right),$
where $\overline{x}=\frac{1}{N}\sum_{k=1}^{N}x^{k}$.
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.15) we find
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$
$\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}V_{\psi}(x^{*},x^{1})+2\sqrt{N}\right)$
$=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.3) we find
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$
$\leq\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
Remark 4.6.
By setting any $m>-1$ in (4.3) (in Theorem 4.3), with the assumption that $h(x^{1})=0$, the convergence rate of Algorithm 2 with the time-varying step sizes given in (3.6), will be optimal without the presence of a $\log(N)$ factor, namely $O\left(\frac{1}{\sqrt{N}}\right)$. In comparison with the sub-optimal convergence rate (4.12), when $m>0$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points.
5. Numerical experiments
In process. The goal of the experiments is to show the effects of the proposed weighting scheme in the convergence of the considered algorithms in this paper, for some problems with a geometrical nature.